# Characterization of Gaussian Tensor Ensembles

### v1

@@ 1.3 Gaussian Tensor Ensembles
Gaussian Orthogonal Tensor Ensemble.
A real symmetric tensor $H\in\mathcal{S}_{p}(N)$ belongs to the $\textit{GOTE}(0,1)$ if, as a tensor-valued random variable, it has a density with respect to the natural Lebesgue measure on $\mathcal{S}_{p}(N)$ proportional to :
@@
The law of such a tensor is orthogonally invariant because the density $f$ only depends on the Frobenius norm, and for all $U\in\mathrm{O}(N),\ \ \|U\cdot H\|_{\mathrm{F}}=\|H\|_{\mathrm{F}}$. Moreover, taking into account the symmetry, this also implies that
$H_{i_{1},\ldots,i_{p}}\sim\mathcal{N}(0,\frac{p}{\Gamma_{i_{1},\ldots,i_{p}}}).$
Note that one could also write the variance as
$\sigma_{i_{1},\ldots,i_{p}}^{2}:=\frac{p}{\Gamma_{i_{1},\ldots,i_{p}}}=\frac{1}{(p-1)!}\prod_{j=1}^{N}c_{j}(i_{1},\ldots,i_{p})!,$
with again $c_{j}(i_{1},\ldots,i_{p})$ the number of occurrences of $j$ in $(i_{1},\ldots,i_{p})$. More generally, we say that $H\in\mathcal{S}_{p}(N)$ belongs to the $\textit{GOTE}(\beta,\gamma)$ if the density is proportional to
@@
Gaussian Unitary Tensor Ensemble.
An hermitian tensor $H\in\mathcal{H}_{p}(N)$ belongs to the $\textit{GUTE}(0,1)$ if, as a tensor-valued random variable, it has a density with respect to the natural Lebesgue measure on $\mathcal{H}_{p}(N)$ proportional to :
@@
The law of such a tensor is unitary invariant, and the symmetries in the hermitian character of the tensor gives that if $(i_{1},\ldots,i_{p})\in\mathfrak{P}$
$H_{i_{1},\ldots,i_{p}}\sim\mathcal{N}(0,\frac{p}{2\Gamma_{i_{1},\ldots,i_{p}}}),$
@@
$H_{i_{1},\ldots,i_{p}}\sim\mathcal{N}(0,\frac{p}{2\Gamma_{i_{1},\ldots,i_{p}}})+i\mathcal{N}^{\prime}(0,\frac{p}{2\Gamma_{i_{1},\ldots,i_{p}}}),$
where $\mathcal{N}$ and $\mathcal{N}^{\prime}$ are two independent Gaussian laws on $\mathbb{R}$. More generally, we say that $H\in\mathcal{S}_{p}(N)$ belongs to the $\textit{GUTE}(\beta,\gamma)$ if the density is proportional to
Remark 3.
One could now define a complex Wigner tensor as $W=H/N^{\frac{p-1}{2}}$ with $H$ an Hermitian tensor whose entries are centered, have finite moments and the same variance as a tensor from the GUTE$(0,1)$, and prove the convergence of the tensor moments as done in our previous works for the real case [1, 2].
Gaussian Symplectic Tensor Ensemble.
Finally, a self-dual hermitian tensor $H\in\mathcal{Q}_{p}(2N)$ belongs to the $\textit{GSTE}(0,1)$ if, as a tensor-valued random variable, it has a density with respect to the natural Lebesgue measure on $\mathcal{Q}_{p}(2N)$ proportional to :
$f(H)\propto\exp{\left(-\frac{2\|H\|_{\mathrm{F}}^{2}}{p}\right)}.$
The law of such a tensor is symplectic invariant. Remark that for a self-dual hermitian tensor the entries of all the $Q^{(\epsilon)}$ are in particular real, so the entries $q:=Q_{i_{1},\ldots,i_{p}}$ of the tensor $Q$ satisfy for $s\in\{1,\ldots,p/2\}$ and $\iota=(\iota_{1},\ldots,\iota_{p})\in\{1,2\}^{p}$,
$q_{\iota}=\bar{q}_{\tau_{s}(\iota)}\text{ if }\iota_{2s-1}=\iota_{2s}\quad\mathrm{and}\quad q_{\iota}=-\bar{q}_{\tau_{s}(\iota)}\text{ if }\iota_{2s-1}\neq\iota_{2s},$
where $\tau_{s}:(\iota_{1},\ldots,\iota_{p})\mapsto(\iota_{1},\ldots,3-\iota_{2s-1},3-\iota_{2s},\ldots,\iota_{p})$ is the partial symmetry. It is similar to what happens in the matrix case where the quaternionic elements, denoted $\mathbb{H}$, are the ones with real coordinates in the quaternion basis $(e_{k})_{0\leq k\leq 3}$, that are matrices of the form $\begin{pmatrix}z&-\omega\\ \bar{\omega}&\bar{z}\end{pmatrix}$. The symmetries in the self-dual hermitian character give this time : if $(i_{1},\ldots,i_{p})\in\mathfrak{P}$,
$q_{1,\ldots,1}\sim\mathcal{N}(0,\frac{p}{4\Gamma_{i_{1},\ldots,i_{p}}})\quad\mathrm{and}\quad q_{1,\ldots,1,2}=0,$
and otherwise, if the indices are not paired,
$q_{1,\ldots,1}\sim\mathcal{N}(0,\frac{p}{4\Gamma_{i_{1},\ldots,i_{p}}})+i\mathcal{N}^{\prime}(0,\frac{p}{4\Gamma_{i_{1},\ldots,i_{p}}})\quad\mathrm{and}\quad q_{1,\ldots,1,2}=\mathcal{N}(0,\frac{p}{4\Gamma_{i_{1},\ldots,i_{p}}})+i\mathcal{N}^{\prime}(0,\frac{p}{4\Gamma_{i_{1},\ldots,i_{p}}}),$
where $\mathcal{N}$ and $\mathcal{N}^{\prime}$ are again independent Gaussian laws on $\mathbb{R}$. More generally, we say that $H\in\mathcal{S}_{p}(N)$ belongs to the $\textit{GSTE}(\beta,\gamma)$ if the density is proportional to
@@ 1.4 Trace invariants
$(H\star_{r}H^{\prime})_{i_{1},\ldots,i_{p+p^{\prime}-2r}}=\sum_{j_{1},\ldots,j_{r}}H_{j_{1},\ldots,j_{r},i_{1},\ldots,i_{p-r}}H^{\prime}_{j_{1},\ldots,j_{r},i_{p-r+1},\ldots,i_{p+p^{\prime}-2r}}.$
To contract along legs different from the first ones, one could previously act on $H$ and $H^{\prime}$ by permutations, and in particular, one may get the tensor resulting of the contraction of the $r$-th leg of $H$ with the $t$-th one of $H^{\prime}$ by $H^{(1,r)}\star_{1}H^{\prime(1,t)}$. The contraction operation is crucial in the tensor setting. In particular, it enables to construct a basis of all the polynomials of the tensor invariant under orthogonal, unitary or symplectic transformation. This basis is given by the so-called trace invariants, see Section 3 for their detailed definition and the proofs that they form a basis in each case. They are encoded by $p$-regular multigraph; where the odd halfedges must be matched with even halfedges for the unitary and symplectic cases. We call the rank of a trace invariant the number of vertices in the associated multigraph.
Rank one and two.
If $p$ is odd, there is no trace invariant of rank one. Otherwise, if $p$ is even, there is only one trace invariant of rank one which is associated to the bouquet graph (one vertex and $p/2$ self-loops, see Figure 1). The rank-one trace invariant is given by the paired trace of $H$ defined as follows :
$\mathrm{Tr}^{\bullet\bullet}(H)=\left\{\begin{array}[]{ll}0&\mbox{if }p\mbox{ odd,}\\ \sum_{i_{1},\ldots,i_{p/2}}H_{i_{1},i_{1},\ldots,i_{p/2},i_{p/2}}&\mbox{if }p\mbox{ even.}\end{array}\right.$
For the case of rank two, the trace invariants are associated to the graphs with two vertices, a number $1\leq r\leq p$ of edges between these two vertices with $p-r$ even, and the other edges are self-loops. A distinguished trace invariant of rank two is particularly interesting : the Frobenius norm $\|H\|_{\mathrm{F}}^{2}$, associated to the melon multigraph, see Figure 1. It corresponds to the case where the two vertices are linked by $r=p$ edges and there is no self-loop; in the hermitian case the $t$-th halfedge of the first vertex is matched with the $(t+1)$-th one of the second vertex ($p+1=1$) and in the self-dual hermitian case the $2t$-th one of a vertex is matched with the $12t-$ 1-th one of the other vertex. The computation proving that the trace invariant associated to the melon is the Frobenius norm is done in Section 3.4. Note finally that we have
$\|H-\beta\mathcal{I}\|_{\mathrm{F}}^{2}/\gamma=a\|H\|_{\mathrm{F}}^{2}+b\mathrm{Tr}^{\bullet\bullet}(H)+c,$
@@
Figure 1: Bouquet graph ($p=10$), melon graph ($p=5$), another trace invariant of rank $2$ ($p=5$), respectively $m_{G}(H)=\mathrm{Tr}^{\bullet\bullet}(H)$, $m_{G}(H)=\|H\|^{2}$, $m_{G}(H)=\sum_{a,b,c,d,e}H_{aabcd}H_{ebcde}$.
Remark 4.
In the vector case, there exists only one trace invariant (associated to the melon graph with two vertices and one edge) given by $\|H\|_{\mathrm{F}}^{2}=|H|^{2}_{2}$. In the matrix case, since the $2$-regular multigraph are exactly the cycles, the trace invariants are the traces of powers of $H$ (it is known since Weyl [19] that $(\mathrm{Tr}(H^{k}))_{k\leq N}$ is a basis of invariants).
@@ 1.5 Maxwell Theorem
The original Maxwell theorem states that a random vector of dimension two or more has independent entries and is rotationally invariant if and only it has a law with a density of the form
$H\mapsto\exp(-a|H|_{2}^{2}+c),$
with $(a,c)\in\mathbb{R}^{*}_{+}\times\mathbb{R}$. The generalization of this result to the order two gives that a real symmetric (resp. hermitian, resp. self-dual hermitian) random matrix of dimension two or more has in the same time independent entries (up to symmetries) and a law invariant by orthogonal (resp. unitary, resp. symplectic) transformation if and only if its law is proportional to the one of the GOE (resp. GUE, resp. GSE) eventually dilated and translated, that is a law with a density of the form
$H\mapsto\exp\left(-a\mathrm{Tr}(H^{2})+b\mathrm{Tr}(H)+c\right),$
@@
Theorem (Main).
For any $p\geq 1$ and $N\geq 2$, a real symmetric (resp. hermitian, resp. self-dual hermitian) random tensor $H$ of order $p$ and dimension $N$ has in the same time independent entries (up to symmetries) and a law invariant by orthogonal (resp. unitary, resp. symplectic) transformation if and only if it belongs to the Gaussian Orthogonal (resp. Unitary, resp. Symplectic) Tensor Ensemble, that is it has a law with a density of the form
@@
$H\mapsto\exp\left(-a\|H\|_{\mathrm{F}}^{2}+b\mathrm{Tr}^{\bullet\bullet}(H)+c\right),$
@@
It is very interesting to understand what does this result means in terms of invariants. The theorem proves that independence of the entries and invariance force the law to depend in the same way on only two very particular trace invariants : the Frobenius norm (associated to the melon graph) and the paired trace (associated to the bouquet graph, existing only when $p$ is even).
Letac extension
There exists an extension of the original Maxwell Theorem due to Letac [11] for $N\geq 3$, with counterexamples for $N=2$. He observes that a random vector $H\in\mathbb{R}^{N}$ has independent entries and is ”isotropic”, that is $\mathbb{P}(H=0)=0$ and $H/|H|_{2}$ is uniformly distributed on the sphere, if and only if its components are independent centered Gaussian variables with the same positive variance. It implies Maxwell Theorem when $N\geq 3$ because rotational invariance and independence of the entries imply isotropy for three or more variables, but isotropy contrasts strongly with rotational invariance for two independent random variables. To the best of the author’s knowledge, no generalization to matrices or higher orders exists, we discuss about it in Section 2.3. We will therefore consider a real symmetric random tensor $H$ with independent entries up to symmetries such that $H/\|H\|_{\mathrm{F}}$ is uniformly distributed on the Frobenius sphere.
1.6 Acknowledgments

### v2

@@ 1.3 Gaussian tensor ensembles
Gaussian orthogonal tensor ensemble. A real symmetric tensor $H\in\mathcal{S}_{p}(N)$ belongs to the $\textrm{GOTE}(0,1)$ if, as a tensor-valued random variable, it has a density with respect to the natural Lebesgue measure on $\mathcal{S}_{p}(N)$ proportional to
@@
The law of such a tensor is orthogonally invariant because the density $f$ only depends on the Frobenius norm, and for all $U\in\mathrm{O}(N)$, $\|U\cdot H\|_{\mathrm{F}}=\|H\|_{\mathrm{F}}$. Moreover, taking into account the symmetry, this also implies that $H_{i_{1},\dots,i_{p}}\sim\mathcal{N}\big(0,\frac{p}{\Gamma_{i_{1},\dots,i_{p}}}\big)$. Note that one could also write the variance as
$\sigma_{i_{1},\dots,i_{p}}^{2}:=\frac{p}{\Gamma_{i_{1},\dots,i_{p}}}=\frac{1}{(p-1)!}\prod_{j=1}^{N}c_{j}(i_{1},\dots,i_{p})!,$
with again $c_{j}(i_{1},\dots,i_{p})$ the number of occurrences of $j$ in $(i_{1},\dots,i_{p})$. More generally, we say that $H\in\mathcal{S}_{p}(N)$ belongs to the $\textrm{GOTE}(\beta,\gamma)$ if the density is proportional to
@@
Gaussian unitary tensor ensemble. An hermitian tensor $H\in\mathcal{H}_{p}(N)$ belongs to the $\textrm{GUTE}(0,1)$ if, as a tensor-valued random variable, it has a density with respect to the natural Lebesgue measure on $\mathcal{H}_{p}(N)$ proportional to
@@
The law of such a tensor is unitary invariant, and the symmetries in the hermitian character of the tensor gives that if $(i_{1},\dots,i_{p})\in\mathfrak{P}$ $H_{i_{1},\dots,i_{p}}\sim\mathcal{N}\big(0,\frac{p}{2\Gamma_{i_{1},\dots,i_{p}}}\big)$, and otherwise, if the indices are not paired,
$H_{i_{1},\dots,i_{p}}\sim\mathcal{N}\left(0,\frac{p}{2\Gamma_{i_{1},\dots,i_{p}}}\right)+{\rm i}\mathcal{N}^{\prime}\left(0,\frac{p}{2\Gamma_{i_{1},\dots,i_{p}}}\right),$
where $\mathcal{N}$ and $\mathcal{N}^{\prime}$ are two independent Gaussian laws on $\mathbb{R}$. More generally, we say that $H\in\mathcal{S}_{p}(N)$} belongs to the $\textrm{GUTE}(\beta,\gamma)$ if the density is proportional to
Remark 1.6.
One could now define a complex Wigner tensor as $W=H/N^{\frac{p-1}{2}}$ with $H$ an Hermitian tensor whose entries are centered, have finite moments and the same variance as a tensor from the GUTE(0,1)(0,1), and prove the convergence of the tensor moments as done in our previous works for the real case [1, 2].
Gaussian symplectic tensor ensemble. Finally, a self-dual hermitian tensor $H\in\mathcal{Q}_{p}(2N)$ belongs to the $\textrm{GSTE}(0,1)$ if, as a tensor-valued random variable, it has a density with respect to the natural Lebesgue measure on $\mathcal{Q}_{p}(2N)$ proportional to
$f(H)\propto\exp{\left(-\frac{2\|H\|_{\mathrm{F}}^{2}}{p}\right)}.$
The law of such a tensor is symplectic invariant. Remark that for a self-dual hermitian tensor the entries of all the $Q^{(\epsilon)}$ are in particular real, so the entries $q:=Q_{i_{1},\dots,i_{p}}$ of the tensor $Q$ satisfy for $s\in\{1,\dots,p/2\}$ and $\iota=(\iota_{1},\dots,\iota_{p})\in\{1,2\}^{p}$}, $q_{\iota}=\bar{q}_{\tau_{s}(\iota)}\text{ if }\iota_{2s-1}=\iota_{2s}$ and $q_{\iota}=-\bar{q}_{\tau_{s}(\iota)}$} if $\iota_{2s-1}\neq\iota_{2s}$}, where $\tau_{s}\colon(\iota_{1},\dots,\iota_{p})\mapsto(\iota_{1},\dots,3-\iota_{2s-1},3-\iota_{2s},\dots,\iota_{p})$ is the partial symmetry. It is similar to what happens in the matrix case where the quaternionic elements, denoted $\mathbb{H}$, are the ones with real coordinates in the quaternion basis $(e_{k})_{0\leq k\leq 3}$, that are matrices of the form $\bigl(\begin{smallmatrix}z&-\omega\\ \bar{\omega}&\bar{z}\end{smallmatrix}\bigr)$}. The symmetries in the self-dual hermitian character give this time: if $(i_{1},\dots,i_{p})\in\mathfrak{P}$},
[…]
@@
$q_{1,\dots,1}\sim\mathcal{N}\left(0,\frac{p}{4\Gamma_{i_{1},\dots,i_{p}}}\right)+{\rm i}\mathcal{N}^{\prime}\left(0,\frac{p}{4\Gamma_{i_{1},\dots,i_{p}}}\right),$
$q_{1,\dots,1,2}=\mathcal{N}\left(0,\frac{p}{4\Gamma_{i_{1},\dots,i_{p}}}\right)+{\rm i}\mathcal{N}^{\prime}\left(0,\frac{p}{4\Gamma_{i_{1},\dots,i_{p}}}\right),$
where $\mathcal{N}$ and $\mathcal{N}^{\prime}$ are again independent Gaussian laws on $\mathbb{R}$. More generally, we say that $H\in\mathcal{S}_{p}(N)$ belongs to the $\textrm{GSTE}(\beta,\gamma)$ if the density is proportional to
@@ 1.4 Trace invariants
$\bigl(H\star_{r}H^{\prime}\bigr)_{i_{1},\dots,i_{p+p^{\prime}-2r}}=\sum_{j_{1},\dots,j_{r}}H_{j_{1},\dots,j_{r},i_{1},\dots,i_{p-r}}H^{\prime}_{j_{1},\dots,j_{r},i_{p-r+1},\dots,i_{p+p^{\prime}-2r}}.$
To contract along legs different from the first ones, one could previously act on $H$ and $H^{\prime}$ by permutations, and in particular, one may get the tensor resulting of the contraction of the $r$-th leg of $H$ with the $t$-th one of $H^{\prime}$ by $H^{(1,r)}\star_{1}H^{\prime(1,t)}$. The contraction operation is crucial in the tensor setting. In particular, it enables to construct a family that generates all the polynomials of the tensor invariant under orthogonal, unitary or symplectic transformation. This complete family is given by the so-called trace invariants, see Section 3 for their detailed definition and the proofs that they form a complete family in each case. They are encoded by $p$-regular multigraph; where the odd halfedges must be matched with even halfedges for the unitary and symplectic cases. More precisely, given a $p$-regular multigraph with $k$ vertices (and $kp/2$ edges), we decorate the vertices by the tensor $H$ and the edges by indices $i_{1},\dots,i_{kp/2}$. Then the trace invariant is given by the sum, over each index going from $1$ to the dimension $N$, of the product over the vertices of the entry of $H$ given by the $p$ indices around the vertex. A detailed example is given in Example 3.1 in the beginning of Section 3 where trace invariants are properly introduced. We call the rank of a trace invariant the number of vertices in the associated multigraph.
Rank one and two. If $p$ is odd, there is no trace invariant of rank one. Otherwise, if $p$ is even, there is only one trace invariant of rank one which is associated to the bouquet graph (one vertex and $p/2$ self-loops, see Figure 1). The rank-one trace invariant is given by the paired trace of $H$ defined as follows:
$\operatorname{Tr}^{\bullet\bullet}(H)=\begin{cases}0&\mbox{if }p\mbox{ odd,}\\ \displaystyle\sum_{i_{1},\dots,i_{p/2}}H_{i_{1},i_{1},\dots,i_{p/2},i_{p/2}}&\text{if }p\mbox{ even.}\end{cases}$
For the case of rank two, the trace invariants are associated to the graphs with two vertices, a number $1\leq r\leq p$ of edges between these two vertices with prp-r even, and the other edges are self-loops. A distinguished trace invariant of rank two is particularly interesting: the Frobenius norm $\|H\|_{\mathrm{F}}^{2}$, associated to the melon multigraph, see Figure 1. It corresponds to the case where the two vertices are linked by $r=p$ edges and there is no self-loop; in the hermitian case the $t$-th halfedge of the first vertex is matched with the $(t+1)$-th one of the second vertex ($p+1=1$) and in the self-dual hermitian case the $2t$-th one of a vertex is matched with the (2t1)(2t-1)-th one of the other vertex. The computation proving that the trace invariant associated to the melon is the Frobenius norm is done in Section 3.4. Note finally that we have
$\|H-\beta\mathcal{I}\|_{\mathrm{F}}^{2}/\gamma=a\|H\|_{\mathrm{F}}^{2}+b\operatorname{Tr}^{\bullet\bullet}(H)+c,$
@@
Figure 1: Bouquet graph ($p=10$), melon graph ($p=5$), another trace invariant of rank 22 ($p=5$), respectively $m_{G}(H)=\operatorname{Tr}^{\bullet\bullet}(H)$, $m_{G}(H)=\|H\|^{2}_{F}$, $m_{G}(H)=\sum_{a,b,c,d,e}H_{aabcd}H_{ebcde}$.
Remark 1.7.
In the vector case, there exists only one trace invariant (associated to the melon graph with two vertices and one edge) given by $\|H\|_{\mathrm{F}}^{2}=|H|^{2}_{2}$. In the matrix case, since the $2$-regular multigraph are exactly the cycles, the trace invariants are the traces of powers of $H$ (it is known since Weyl [19] that $\bigl(\operatorname{Tr}\bigl(H^{k}\bigr)\bigr)_{k\leq N}$ is a basis of invariants).
@@ 1.5 Maxwell theorem
The original Maxwell theorem states that a random vector of dimension two or more has independent entries and is rotationally invariant if and only it has a law with a density of the form $H\mapsto\exp\bigl(-a|H|_{2}^{2}+c\bigr)$}, with $(a,c)\in\mathbb{R}^{*}_{+}\times\mathbb{R}$. The generalization of this result to the order two gives that a real symmetric (resp. hermitian, resp. self-dual hermitian) random matrix of dimension two or more has in the same time independent entries (up to symmetries) and a law invariant by orthogonal (resp. unitary, resp. symplectic) transformation if and only if its law is proportional to the one of the GOE (resp. GUE, resp. GSE) eventually dilated and translated, that is, a law with a density of the form
$H\mapsto\exp\left(-a\operatorname{Tr}\bigl(H^{2}\bigr)+b\operatorname{Tr}(H)+c\right),$
@@
Theorem 1.8 (main).
For any $p\geq 1$ and $N\geq 2$, a real symmetric ((resp. hermitian, resp. self-dual hermitian)) random tensor $H$ of order $p$ and dimension $N$ has in the same time independent entries ((up to symmetries)) and a law invariant by orthogonal ((resp. unitary, resp. symplectic)) transformation if and only if it belongs to the Gaussian orthogonal ((resp. unitary, resp. symplectic)) tensor ensemble, that is, it has a law with a density of the form
@@
$H\mapsto\exp\left(-a\|H\|_{\mathrm{F}}^{2}+b\operatorname{Tr}^{\bullet\bullet}(H)+c\right),$
@@
It is very insightful to understand what does this result means in terms of invariants. The theorem proves that independence of the entries and invariance force the law to depend in the same way on only two very particular trace invariants: the Frobenius norm (associated to the melon graph) and the paired trace (associated to the bouquet graph, existing only when $p$ is even).
Letac extension. There exists an extension of the original Maxwell theorem due to Letac [11] for $N\geq 3$, with counterexamples for $N=2$. He observes that a random vector $H\in\mathbb{R}^{N}$ has independent entries and is “isotropic”, that is, $\mathbb{P}(H=0)=0$ and $H/|H|_{2}$ is uniformly distributed on the sphere, if and only if its components are independent centered Gaussian variables with the same positive variance. It implies Maxwell theorem when $N\geq 3$ because rotational invariance and independence of the entries imply isotropy for three or more variables, but isotropy contrasts strongly with rotational invariance for two independent random variables. To the best of the author’s knowledge, no generalization to matrices or higher orders exists, we discuss about it in Section 2.3. We will therefore consider a real symmetric random tensor $H$ with independent entries up to symmetries such that $H/\|H\|_{\mathrm{F}}$ is uniformly distributed on the Frobenius sphere.